OpenStudy (anonymous):

HELP PLEASE Describe the differences between the graph of y = –3(x + 7)2 – 10 and the standard position graph of y = x2.

OpenStudy (goten77):

let me retype this i just did it all wrong <.<

OpenStudy (goten77):

\[-3(x+7)^{2}-10=-3(x+7)(x+7)- 10 = -3(x ^{2}+14x + 49)-10 = -3x ^{2} -42x -157\]

OpenStudy (goten77):

hmm that last part should show \[-3x ^{2} -42x - 157\]

OpenStudy (goten77):

k and when we graph the 2nd 1 we gotta zoom out since its y intercept is -157 and heres what the 2 graps look like when compared by drawing